# Deep Network Pruning: A Comparative Study on CNNs in Face Recognition

### v1

@@ 4 Results and Discussion
We apply pruning to the networks using a random 25% of the VGGFace2 training set to compute the filter importance scores. We then remove 1% of the lowest-scoring filters after each epoch. The optimizer is SGDM with mini-batch 128 and learning rate 0.01. Figure 4 (blue curves) shows the EER of the pruned networks on VGGFace2-Pose. The baseline EER at $x$=0 (no pruning) is 6.07% (SqueezeNet), 3.75% (MobileNetv2) and 3.84% (ResNet50) in one-to-one comparisons (single-image template), and 0.52%, 0.14% and 0.13% for five-to-five comparisons. The deeper MobileNetv2 and ResNet50 outperform the lighter SqueezeNet, with MobileNetv2 matching or exceeding ResNet50 with fewer parameters and size, demonstrating the effectiveness of the MobileNetv2 design to provide good accuracy with a reduced network.
@@
To regain accuracy losses, pruned networks are retrained again on VGGFace2. The results are given in Figure 4 as well. We use a starting learning rate of 0.01 (red curve) and 0.001 (orange). We hypothesise that, since the network has been already trained on the same database, starting with a high rate may be counterproductive. However, this is not the case, since the best results are with a starting rate of 0.01. All architectures maintain accuracy up to a certain pruning percentage. In addition, using five images per template allows higher pruning sparsities. Combining several images may produce that the most relevant features of each identity are kept, while random information is averaged out [Caldeira et al., 2024]. However, similarly as above, the behaviour per network differs. SqueezeNet maintains performance until 15-20% (one-to-one) or 20-30% pruning (five-to-five). ResNet remains stable even with high pruning, and the performance gap w.r.t. the unpruned network is eliminated. On the other hand, the sensitiveness of MobileNetv2 to pruning is substantially reduced after retraining, maintaining a stable performance until 20-30% pruning in both one-to-one and five-to-five comparisons. At 40% or 50% pruning, it remains competitive, and on par or even better than the bigger ResNet50. Indeed, the EERs of MobileNetv2 are consistently smaller than ResNet50 (especially in one-to-one comparisons), suggesting that MobileNetv2 is a very robust architecture with capabilities comparable to much larger networks.
@@
We analyze pruning effect on the networks themselves by plotting learnables and embedding dimensionality (Figure 5). In all networks, learnables decrease slowly until 20-30% of the filters are removed, then the decrease is sharper, so the filters that are removed first do not affect a high amount of channels. There is a small exception with SqueezeNet, which has a pronounced drop between 10-15%, but the decrease stabilizes afterwards. Embedding size remains constant until certain pruning levels which, again, is different per network. ResNet50 starts reducing its output dimensionality with less than 10% pruning, SqueezeNet at $\sim$20%, and MobileNetv2 at $\sim$40%. A reduction in the embedding indicates that pruning affects the last layer. Such layer is supposed to be the more specific for the task, so it could be expected that it is not affected in the early pruning stages [Caldeira et al., 2024]. However, this does not hold with ResNet50. Nevertheless, ResNet50 is substantially bigger (Table 1), its embedding duplicates others (2048 vs. 1280 or 1000), but its performance is not substantially better (Figure 4). Research has shown that feature spaces of 512 or 1024 contain sufficient degrees of freedom to encode a face, allowing further reduction with minor accuracy drops [O’Toole et al., 2018]. Thus, it is reasonable that the large ResNet50 embedding is affected from the beginning.
@@
Fig. 8: Effect of pruning in the amount of filters per layer. For each layer, the y-axis indicate the amount of filters between 0% and 20% pruning. The circles indicate the layers with the fastest reduction in the amount of filters. The layer names correspond to the networks’ release of Matlab r2023b.
@@
The latter observations suggest that ResNet50 is over-dimensionated on several aspects. At 71% pruning, it has ten times more parameters than MobileNetv2 (6.97M vs. 0.67M), but not significantly better accuracy: 4.83% vs. 5.29% in one-to-one comparisons (8.7% decrease), or 0.15% vs. 0.31% in five-to-five comparisons (51.6% decrease). Indeed, at 99% pruning, ResNet50 learnables are comparable to MobileNetv2 without any pruning. To better analyze this effect, we plot in Figure 6 the accuracy against the number of learnables. In one-to-one comparisons, when ResNet50 and MobileNetv2 have similar parameters ($\sim$1.07M), MobileNetv2 has a 2% smaller EER. This occurs when MobileNetv2 is pruned at 26% and ResNet50 at 99%. Despite similar depth (conv layers) and parameters, MobiletNetv2 seems fundamentally better. Figure 7 shows the filters per layer of MobileNetv2 at 26% pruning and ResNet50 at 99%. ResNet50 layers mostly have less than 150 filters, even in the last stage of bottleneck layers originally designed to increase dimensionality by 4$\times$ (those containing ‘branch2c’ in the name). Also, the last part of the network has 250 filters, which contrasts to the 2048 of the unpruned network. On the other hand, MobileNetv2 maintains a more prominent difference inside bottlenecks, Also, filter count increases substantially towards the end of the network, when the layers are supposed to be more specific for the task, which may explain its superior performance w.r.t. ResNet50.
@@
Interestingly, in the 1-1.1M parameters range, ResNet50 (99% pruned) and SqueezeNet ($<$10% pruned) perform similarly, despite the simpler structure of SqueezeNet. However, the smaller embedding size of ResNet50 (245 elements, Figure 5) vs. SqueezeNet (1000) suggests that the most intrincate structure of ResNet can provide similar performance with a smaller embedding. On the other hand, differences between ResNet50 and MobileNetv2 disappear with five images per template (bottom plot of Figure 6), compensating for the observed limitations of ResNet50. It can be also seen that the gap of SqueezeNet remains in five-to-five comparisons, showing the limitations of such a simpler architecture.
@@
We further analyze the effect of pruning on the networks by plotting (Figure 8) the filters per layer for different pruning degrees (up to 20%). Filters removed first (marked with black circles) belong to layers with the most channels, which are dedicated to dimensionality increase. Similarly to the embedding size, these high-dimensional spaces seem over-dimensioned for the task. This is obviously more prominent in the network with the biggest dimensional spaces (ResNet50). At 20% pruning, ResNet50 layers of original dimensionality 2048, 1024, and 512 are reduced to less than 1000, 540, and 255 channels respectively (less than half). This aligns with previous findings (Figure 7) showing reduced dimensionality differences inside ResNet50 bottlenecks throughout pruning. Thus, it makes sense that high-dimensional channels are affected the most.
@@ 5 Conclusions
Popular face recognition technologies use large Convolutional Neural Networks (CNNs) [Sundararajan and Woodard, 2018] with hundreds of megabytes and/or millions of parameters, which can be impractical for low-capacity hardware. We are thus interested in methods to achieve lighter CNN architectures without sacrificing performance. Here, we explore the application of a network pruning method based on Taylor scores which only require the backpropagation gradient to measure the impact of filters on the classification error. This allows iterative removal of least contributing filters. The method is evaluated on three architectures with varying parameters and depth. Each network allows a different degree of pruning without substantial performance loss, based on its initial size. For example, the heaviest architecture allows up to 99% pruning, while the second heaviest allows below 30%. However, at these points, both have similar learnables ($\sim$1M) and performance. Thus, larger networks are more over-dimensioned (at least for the architectures evaluated here). We also observed that filters with the most channels are removed first, regardless of the architecture. This suggests that the high dimensional spaces of the three CNNs employed seem to be highly over-dimensionated for the task at hand. Furthermore, we found that using multiple input images to model identity of users strengthens the network, allowing higher pruning sparsities.
In future work, we are looking at other alternatives for network compression to further optimize the balance between network size and performance, such as pruning combined with Knowledge Distillation (KD). For example, the pruned networks could be trained to mimic the embedding of a more powerful one, e.g. [Deng et al., 2019a]. However, this would require that both the teacher and student networks have the same embedding size, which could be achieved by blocking pruning in the last convolutional layer. It has also been suggested that feature spaces of CNN trained in the same task can be linearly mapped [McNeely-White et al., 2020], which would solve the issue of embeddings of different size. The combination of several compression techniques, including quantization, KD and pruning, have also shown to be benefecial to achieve model reductions [Almadan and Rattani, 2023], and it will also be the source of future research.
@@ Acknowledgments
This work was partly done while F. A.-F. was a visiting researcher at the University of the Balearic Islands. F. A.-F., K. H.-D., and J. B. thank the Swedish Research Council (VR) and the Swedish Innovation Agency (VINNOVA) for funding their research. Author J. M. B. thanks the project EXPLAINING - ”Project EXPLainable Artificial INtelligence systems for health and well-beING”, under Spanish national projects funding (PID2019-104829RA-I00/AEI/10.13039/501100011033).

### v2

@@ 4 Results and Discussion
We apply pruning to the networks using a random 25% of the VGGFace2 training set to compute the filter importance scores. We then remove 1% of the lowest-scoring filters after each epoch. The optimizer is SGDM with mini-batch 128 and a learning rate of 0.01. Figure 4 (blue curves) shows the EER of the pruned networks on VGGFace2-Pose. The baseline EER at $x$=0 (no pruning) is 6.07% (SqueezeNet), 3.75% (MobileNetv2) and 3.84% (ResNet50) in one-to-one comparisons (single-image template), and 0.52%, 0.14% and 0.13% for five-to-five comparisons. The deeper MobileNetv2 and ResNet50 outperform the lighter SqueezeNet, with MobileNetv2 matching or exceeding ResNet50 with fewer parameters and size, demonstrating the effectiveness of the MobileNetv2 design to provide good accuracy with a reduced network.
@@
To regain accuracy losses, pruned networks are retrained again on VGGFace2. The results are given in Figure 4 as well. We use a starting learning rate of 0.01 (red curve) and 0.001 (orange). We hypothesise that since the network has already been trained on the same database, starting with a high rate may be counterproductive. However, this is not the case, since the best results are with a starting rate of 0.01. All architectures maintain accuracy up to a certain pruning percentage. In addition, using five images per template allows higher pruning sparsities. Combining several images may produce that the most relevant features of each identity are kept, while random information is averaged out [Caldeira et al., 2024]. However, similarly to the above, the behaviour per network differs. SqueezeNet maintains performance until 15-20% (one-to-one) or 20-30% pruning (five-to-five). ResNet remains stable even with high pruning, and the performance gap w.r.t. the unpruned network is eliminated. On the other hand, the sensitiveness of MobileNetv2 to pruning is substantially reduced after retraining, maintaining a stable performance until 20-30% pruning in both one-to-one and five-to-five comparisons. At 40% or 50% pruning, it remains competitive, and on par or even better than the bigger ResNet50. Indeed, the EERs of MobileNetv2 are consistently smaller than ResNet50 (especially in one-to-one comparisons), suggesting that MobileNetv2 is a very robust architecture with capabilities comparable to much larger networks.
@@
We analyze the pruning effect on the networks themselves by plotting learnables and embedding dimensionality (Figure 5). In all networks, learnables decrease slowly until 20-30% of the filters are removed, then the decrease is sharper, so the filters that are removed first do not affect a high amount of channels. There is a small exception with SqueezeNet, which has a pronounced drop between 10-15%, but the decrease stabilizes afterwards. Embedding size remains constant until certain pruning levels, which, again, is different per network. ResNet50 starts reducing its output dimensionality with less than 10% pruning, SqueezeNet at $\sim$20%, and MobileNetv2 at $\sim$40%. A reduction in the embedding indicates that pruning affects the last layer. Such layer is supposed to be the more specific for the task, so it could be expected that it is not affected in the early pruning stages [Caldeira et al., 2024]. However, this does not hold with ResNet50. Nevertheless, ResNet50 is substantially bigger (Table 1), its embedding duplicates others (2048 vs. 1280 or 1000), but its performance is not substantially better (Figure 4). Research has shown that feature spaces of 512 or 1024 contain sufficient degrees of freedom to encode a face, allowing further reduction with minor accuracy drops [O’Toole et al., 2018]. Thus, it is reasonable that the large ResNet50 embedding is affected from the beginning.
@@
Fig. 8: Effect of pruning in the amount of filters per layer. For each layer, the y-axis indicates the amount of filters between 0% and 20% pruning. The circles indicate the layers with the fastest reduction in the amount of filters. The layer names correspond to the networks’ release of Matlab r2023b. Better if zoomed.
@@
The latter observations suggest that ResNet50 is over-dimensioned in several aspects. At 71% pruning, it has ten times more parameters than MobileNetv2 (6.97M vs. 0.67M), but not significantly better accuracy: 4.83% vs. 5.29% in one-to-one comparisons (8.7% decrease), or 0.15% vs. 0.31% in five-to-five comparisons (51.6% decrease). Indeed, at 99% pruning, ResNet50 learnables are comparable to MobileNetv2 without any pruning. To better analyze this effect, we plot in Figure 6 the accuracy against the number of learnables. In one-to-one comparisons, when ResNet50 and MobileNetv2 have similar parameters ($\sim$1.07M), MobileNetv2 has a 2% smaller EER. This occurs when MobileNetv2 is pruned at 26% and ResNet50 at 99%. Despite similar depth (conv layers) and parameters, MobiletNetv2 seems fundamentally better. Figure 7 shows the filters per layer of MobileNetv2 at 26% pruning and ResNet50 at 99%. ResNet50 layers mostly have less than 150 filters, even in the last stage of bottleneck layers originally designed to increase dimensionality by 4$\times$ (those containing ‘branch2c’ in the name). Also, the last part of the network has 250 filters, which contrasts with the 2048 of the unpruned network. On the other hand, MobileNetv2 maintains a more prominent difference inside bottlenecks, Also, filter count increases substantially towards the end of the network, when the layers are supposed to be more specific for the task, which may explain its superior performance w.r.t. ResNet50.
@@
Interestingly, in the 1-1.1M parameters range, ResNet50 (99% pruned) and SqueezeNet ($<$10% pruned) perform similarly, despite the simpler structure of SqueezeNet. However, the smaller embedding of ResNet50 (245 elements, Figure 5) vs SqueezeNet (1000) suggests that the most intricate structure of ResNet can provide similar performance with a smaller embedding. On the other hand, differences between ResNet50 and MobileNetv2 disappear with five images per template (bottom plot of Figure 6), compensating for the observed limitations of ResNet50. It can also be seen that the gap of SqueezeNet remains in five-to-five comparisons, showing the limitations of such a simpler architecture.
@@
We further analyze the effect of pruning by plotting (Figure 8) the filters per layer for different pruning degrees (up to 20%). Filters removed first (marked with black circles) belong to layers with the most channels, which are dedicated to dimensionality increase. Similarly to the embedding size, these high-dimensional spaces seem over-dimensioned for the task. This is obviously more prominent in the network with the biggest dimensional spaces (ResNet50). At 20% pruning, ResNet50 layers of original dimensionality 2048, 1024, and 512 are reduced to less than 1000, 540, and 255 channels, respectively (less than half). This aligns with previous findings (Figure 7) showing reduced dimensionality differences inside ResNet50 bottlenecks throughout pruning. Thus, it makes sense that high-dimensional channels are affected the most.
Fig. 9: Average heatmaps of the original (unpruned) and different pruned versions of the networks. The average face of VGGFace2-Pose is also shown for reference.
Fig. 10: Histogram of PSNR between the heatmaps of the original (unpruned) and different pruned versions of the networks. The top left of each subfigure shows the one-to-one EER of the unpruned network, whereas the numbers above each curve are the EER of the pruned counterparts.
Fig. 11: Individual heatmaps of two images of VGGFace2-Pose with the original (unpruned) and different pruned versions of the networks.
Finally, we study the impact of pruning on the heatmaps that quantify the most relevant pixels for the networks. We do so via the Local Interpretable Model-Agnostic Explanations (LIME) as employed in [Alonso-Fernandez et al., 2023a]. Figure 9 depicts the average heatmaps of each network on the VGGFace2-Pose set. We show selected pruning levels (26 and 96%), both with and without retraining. As can be seen, the average heatmaps without retraining (columns 2 and 4) depart progressively from the original map of the unpruned networks (column 1). The effect is more remarkable with MobileNetv2, which was observed to be very sensitive to pruning. After retraining (columns 3, 5), the heatmaps recover to a certain extent the shape seen in the first row, which focuses its attention mostly on the nose and eyes area. We also study the effect at the individual image level by showing in Figure 10 the histograms of the PSNR between the heatmaps of the original (unpruned) network and the different pruned counterparts. As can be expected, retrained networks (solid curves) show higher PSNRs than non-retrained ones (dashed curves). Similarly, pruning at 26% (red curves) shows higher PSNRs than pruning at 96% (green curves). It must be taken into account that at 26 and 96% pruning, the number of parameters and performance of each network was different (Figure 6). Thus, the ranges of PSNR for each network in the x-axis are different, being closely tied to the EER (shown above each curve). Thus, the more the heatmaps depart from the original (unpruned) version, the higher the EER given by such network. Lastly, We give some examples of individual heatmaps (Figure 11) corresponding to the first two users of VGGFace2-Pose. Similar considerations than with the average heatmaps above can be made. At 0% pruning, the networks focus the attention mostly on the nose area and, to a lesser extent, on the eye area. In one example, the networks also focus on the forehead region. The heatmaps at 26 and 96% pruning without retraining (rows 2, 4) depart progressively from the original map of the unpruned networks (row 1). In those cases, the networks sometimes focus their attention on other regions of the image, especially at 96% pruning. After retraining (rows 3, 5), the networks mostly focus their attention again on the same regions as the original map. There are some exceptions, in particular, SqueezeNet and MobileNetv2 at 96% pruning. However, at this working point, these networks show a degraded EER compared to the unpruned network (Figure 10), so we can expect the heatmaps to differ as well. On the other hand, the EER of ResNet50 does not degrade as much comparatively. Therefore, there is more similarity between heatmaps at different pruning levels.
@@ 5 Conclusions
Popular face recognition technologies use large Convolutional Neural Networks (CNNs) [Sundararajan and Woodard, 2018] with hundreds of megabytes and millions of parameters, which can be impractical for low-capacity hardware. We are thus interested in methods to achieve lighter CNN architectures without sacrificing performance. Here, we explore the application of a network pruning method based on Taylor scores, which only requires the backpropagation gradient to measure the impact of filters on classification errors. This allows iterative removal of the least contributing filters. The method is evaluated on three architectures with varying parameters and depth. Each network allows a different degree of pruning without substantial performance loss, based on its initial size. For example, the heaviest architecture allows up to 99% pruning, while the second heaviest allows below 30%. However, at these points, both have similar learnables ($\sim$1M) and performance. Thus, larger networks are more over-dimensioned (at least for the architectures evaluated here). We also observed that filters with the most channels are removed first, regardless of the architecture. This suggests that the high dimensional spaces of the CNNs employed seem to be highly over-dimensioned for the task. Furthermore, we found that using multiple input images to model the identity of users strengthens the network, allowing higher pruning sparsities. In the last stage of our experimental setup, we quantify (via heatmaps) the most relevant pixels for the networks. The heatmaps are observed to progressively change with pruning, focusing on regions different from the original unpruned network. After retraining, they somehow resemble the original heatmaps again to an extent that is dependent on the EER gap between the pruned and unpruned networks.
In future work, we are looking at other alternatives for network compression to further optimize the balance between network size and performance, such as pruning combined with Knowledge Distillation (KD). For example, the pruned networks could be trained to mimic the embedding of a more powerful one, e.g. [Deng et al., 2019]. However, this would require that both the teacher and student networks have the same embedding size. We are working on this from two perspectives: $i$) by blocking pruning in the last convolutional layer, but also $ii$) by first retraining the original CNNs to have the same output size. The latter approach takes considerably longer, but it will allow the student and teacher to use different architectures. It has also been suggested that feature spaces of CNN trained in the same task can be linearly mapped [McNeely-White et al., 2020], which would solve the issue of embeddings of different sizes. The combination of several compression techniques, including quantization, KD and pruning, has also shown to be beneficial to achieve model reductions [Almadan and Rattani, 2023], and it will also be the source of future research. Finally, we are working on extending this study to other face databases of wide use in the literature [Kolf et al., 2023].
@@ Acknowledgments
This work was partly done while F. A.-F. was a visiting researcher at the University of the Balearic Islands. F. A.-F., K. H.-D., and J. B. thank the Swedish Research Council (VR) and the Swedish Innovation Agency (VINNOVA) for funding their research. This work is part of the Project PID2022-136779OB-C32 (PLEISAR) funded by MICIU/ AEI /10.13039/501100011033/ and FEDER, EU